OpenStudy (anonymous):

Yuto and Lian are at train stations 1,880 kilometers apart. Yuto boards a train heading east at an average speed of 220 kilometers per hour. At the same time, Lian boards a train heading west on a parallel track at an average speed of 250 kilometers per hour. How far has Lian traveled when the two trains pass each other? 470 kilometers 880 kilometers 940 kilometers 1,000 kilometers

OpenStudy (anonymous):

@alex1248191

OpenStudy (anonymous):

basically set the equation up like the other questions and solve it

OpenStudy (shinalcantara):

velocity is distance over time \[v(velocity) = \frac{ d(distance) }{ t(time) }\] \[d = vt\] since you're given with the total distance then you'll have the equation 220t + 250t = 1,880 470t =1880 t = 4hours since the train to which Lian is, travels at 250kph then 250t = 250(4) = 1000 km